| 
HOME
ESR
EXPERIMENTS
LABS etc
 | 
| 
E121
E127
E132
E125
E143
laser_cooling_2021
E142
Ê128
E146
E0052
E0018
E0028
E0038
E0075
E0203_E0018
E0052_2025
 | 
|   masses & isomeres 2025, Page 12 of 17  | 
Not logged in | 
  | 
   | 
| New entries since: | Thu Jan  1 01:00:00 1970 |    | 
   Thu Apr  3 03:48:40 2025, JEL, ZN, Runs, Screenshots    
 |  
   Thu Apr  3 02:59:15 2025, JEL, ZN, Runs, Screenshot, beam received, baem lost, beam back again    
 |  
   Thu Apr  3 02:52:27 2025, Zac, Johan, General, Beam is back at 02:52 
 |  
   Thu Apr  3 02:01:29 2025, Zac, Johan, General, Quadrupole issues 
 |  
   Thu Apr  3 01:56:57 2025, Zac, Johan, General, Problems with one of the quadrupoles 
 |  
   Thu Apr  3 01:29:49 2025, Zac, Johan, General, Beam is back at 01:27 
 |  
   Thu Apr  3 01:19:44 2025, Zac, Johan, General, Beam fluctuations 
 |  
   Thu Apr  3 01:14:16 2025, Zac, Johan, General, Beam is back at 01:12 
 |  
   Thu Apr  3 00:49:00 2025, Zac, Johan, General, Beam lost at 00:46 
 |  
   Wed Apr  2 23:40:23 2025, JEL, ZN, General, Screenshots taken just as beam came back    
 |  
   Wed Apr  2 23:33:43 2025, Zac, Accelerator, Beam is gone at 23:29 
 |  
   Wed Apr  2 23:27:35 2025, Zac, Johan, Accelerator, Beam is back 23:25 
 |  
   Wed Apr  2 23:15:52 2025, Zac, Johan, Accelerator, No injections from SIS 
 |  
   Wed Apr  2 22:47:25 2025, ID, HA (Radio Shahab as background), General, Lassie and Schottky TV screentshots    
 |  
   Wed Apr  2 21:22:41 2025, ID, HA, General, Lassie Come-Home Screenshots (wuff!)    
 |  
   Wed Apr  2 20:57:34 2025, Shahab, Detectors, 10GB cables for the DAQ rack 
 |  
   Wed Apr  2 20:15:29 2025, ID, HA, General, Lassie screenshots + Schottky TV    
 |  
   Wed Apr  2 19:13:50 2025, ID, HA, General, Lassie screenshots and Schottky screenshot    
 |  
   Wed Apr  2 18:31:52 2025, ID, General, FYI for morning shifts Thu 
 |  
   Wed Apr  2 18:22:30 2025, ID, General, Spill cycle distribution 
 |   
 |